OpenStudy (anonymous):

-2[(4p + 1) -(3p - 1)] = 5(3 - p) - 9 Show steps please!! :)

OpenStudy (anonymous):

umm this is engish but ok i'll try Epand brakets -2[4p+1-3p+1] =15-5p-9 -8p-2+6p-2=15-5p-9 collect like terms -2p-4=6-5p -2p+5p=+4+6 3p=10 isolate P p=3.3333333...

OpenStudy (trancenova):

Hello, please ask your questions in the correct group - looks like this one accidentally ended up in English - so its now closed.